OpenStudy (anonymous):

AB passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to AB ?

OpenStudy (dogzcatz):

do you know how to solve these kinds of question shut up

OpenStudy (anonymous):

A. 5x − 3y = 0 B. -x + 3y = 0 C. -5x − 3y = 0 D. 3x + 5y = 0 E. -3x + 5y = 0

OpenStudy (dogzcatz):

Do you ever type what you hear I didn't mean shut up at you

OpenStudy (dogzcatz):

either choice A or C which one do you think

OpenStudy (dogzcatz):

I do know which one, but from what you know how would you write the equation

OpenStudy (dogzcatz):

It's C.

OpenStudy (anonymous):

lol okay thank you. sorry I was away from the computer. but thank you means a lot